Nancy perkins borrow $2000 at 5% and she is going to make annual payments of $734.42.How long before nancy off the loan.


1
Expert's answer
2021-05-06T13:29:44-0400

total amount"= \frac{105}{100} \times 2000"

$ 2100

Annual payment= $734

"Duration=\frac{2100}{734}"

= 2.86

Answer 3 years